### v1

# The Rhythm of Aging Stability and Drift in Human Senescence

Abstract
Human aging is marked by a steady rise in mortality risk with age — a process demographers describe as senescence. While life expectancy has improved dramatically over the past century, a fundamental question remains: is the rate at which mortality accelerates biologically fixed, or has it shifted across generations? Vaupel’s hypothesis suggests that the pace of aging is stable — that humans are not aging more slowly, but simply starting later. To test this, we analyze cohort mortality data from France, Denmark, Italy, and Sweden. We use a two-step framework to first isolate senescent mortality, then decompose the Gompertz slope into three parts: a biological constant, a potential trend, and a cumulative period effect. The results show that most variation in the rate of aging is not biological in origin. Once non-senescent deaths and historical shocks are accounted for, the Gompertz slope is remarkably stable. The fluctuations we see are not signs of changing senescence, but echoes of shared history. Aging itself, it seems, has stayed the same. These findings suggest that while longevity has shifted, the fundamental rhythm of human aging may be biologically fixed — shaped not by evolution, but by history.
Keywords: Actuarial senescence, Gompertz law, Rate of aging, Cohort analysis, Period effects
1 Introduction
Aging is the gradual decline in physiological function — what we see as graying hair, slower steps, and growing vulnerability to illness and injury. Beneath these visible signs lies senescence, the biological process that drives aging (Comfort, , 1964). Demographers focus on actuarial senescence — the age-related rise in mortality risk — which, in most adult populations, follows an exponential curve.
This curve can be modeled using the Gompertz law of mortality (Gompertz, , 1825), where the force of mortality increases exponentially with age. The steepness of this curve is captured by a single parameter, $b$, and can be interpreted as the rate at which mortality accelerates. A higher $b$ means mortality rises more steeply with age; a lower $b$ suggests a slower pace of senescence.
At the same time, people around the world are living longer than ever before. Life expectancy has risen steadily for over a century, not just for a few, but across entire populations (Oeppen and Vaupel, , 2002). Since the mid-19th century, best-practice life expectancy has increased by roughly 2.5 years per decade, driven largely by sustained reductions in old-age mortality (Vaupel et al., , 2021). This is one of the great successes of modern societies. But it raises a deeper question: is aging itself changing? Are people aging more slowly — or are they simply starting the aging process later?
Vaupel, (2010) proposed what is now known as Vaupel’s hypothesis: that the rate of aging, $b$, is biologically constant. From this perspective, people are not aging more slowly; they are aging later. The slope of mortality remains the same — it’s just been pushed forward in time. Under this view, gains in life expectancy reflect delayed aging, not a change in the underlying biology of senescence.
The mathematical foundation behind this hypothesis is grounded in the gamma-Gompertz model, where the exponential increase in the hazard of death due to aging is modulated by unobserved individual frailty (Vaupel et al., , 1979; Vaupel and Missov, , 2014). In this framework, even when populations become more heterogeneous, the rate of senescence itself can remain stable.
If the pace of aging is fixed, then aging is not speeding up or slowing down — it’s just being postponed (Vaupel, , 2010). But if the rate of aging is truly changing, then something more fundamental is happening (Kirkwood and Austad, , 2000). It would mean that the biology of aging itself is evolving — or being altered by the environment, behavior, or historical events (Finch and Crimmins, , 2004; Crimmins and Beltrán-Sánchez, , 2011).
Yet when researchers estimate $b$ across populations and birth cohorts, they often find small but persistent variations (Barbi et al., , 2003; Zarulli, , 2013; Salinari and De Santis, , 2014). Some studies reject the hypothesis that $b$ is stable; others observe fluctuations that drift over time (Salinari and De Santis, , 2020; Zarulli et al., , 2012). These patterns raise a deeper question: are we seeing real biological change in how humans age — or are we seeing something else?
One possibility is that what looks like change in $b$ is not biological at all, but historical. Period events — such as World War I, the 1918 flu, or World War II — affect many cohorts at once, just at different ages (Vallin and Meslé, , 2004). These shocks strike in calendar time, but their effects are smeared across birth cohorts. If such events have lasting consequences, they could subtly distort cohort-level mortality patterns — not in sharp jumps, but through slow, cumulative shifts (Horiuchi, , 2003; Zarulli et al., , 2012).
This matters because when we estimate $b$ cohort by cohort, we assume we are tracing a biological process. But we may be picking up the long echo of shared historical event — nudging the estimated rate of aging up or down, year by year, in ways that mimic biological drift. Over time, these nudges can pile up, creating what looks like a change in the slope of mortality — even if aging itself hasn’t changed at all (Zarulli et al., , 2012; Salinari and De Santis, , 2014; Horiuchi and Wilmoth, , 1998).
These kinds of latent effects are hard to see directly. But their signature is familiar: they accumulate gradually, move in one direction for a while, then turn. We can think of this as a stochastic process—more specifically, a random walk (e.g., Grimmett and Stirzaker, , 2020; Hamilton, , 2020). If period-driven shocks behave this way, they could generate the illusion of a shifting $b$ — even when the underlying biology stays constant (Yashin et al., , 2000).
This paper asks: Is the rate of aging truly changing, or is the variation we observe across cohorts the result of cumulative period shocks that mimic change? We approach this question by decomposing the estimated rate of aging into its possible parts: a biological constant, a deterministic trend, a latent accumulation of shared historical effects, and residual noise. This structure allows us to see not just whether $b$ changes — but why it appears to.
2 Background and Motivation
Estimates of the rate of aging, $b$, often vary across birth cohorts. Sometimes these values drift gently up or down; other times they fluctuate without clear direction. Many of these differences are statistically significant, but their biological interpretation remains uncertain. Do these fluctuations reflect real variation in how humans age? Or are they shaped by model assumptions, measurement noise, or shared historical exposures?
James Vaupel’s hypothesis offers a compelling starting point (Vaupel, , 2010). It proposes that the rate at which mortality accelerates with age is a built-in feature of human biology — constant across time and place. This idea is grounded in the gamma-Gompertz model, which accounts for individual frailty while preserving a stable underlying rate of senescence (Vaupel et al., , 1979; Yashin et al., , 2000; Vaupel and Missov, , 2014).
But empirical tests of this hypothesis have yielded mixed findings. Barbi et al., (2003) found that estimates of $b$ for Italian cohorts varied significantly depending on the statistical method used, raising the possibility that apparent changes reflect model sensitivity rather than biological shifts. Similarly, Zarulli et al., (2012); Zarulli, (2013) analyzed the aftermath of large mortality shocks — such as famine and wartime captivity — and found a flattening of the aging rate, likely due to selective survival rather than a true biological response.
Other studies tested the constancy of $b$ more directly. Salinari and De Santis, (2014) rejected the hypothesis that $b$ is stable across countries, sexes, and cohorts, though the differences they observed were modest. A subsequent paper by Salinari and De Santis, (2020) suggested that $b$ might even vary with age, rising before leveling off. However, their model does not separate cohort and age effects, which makes it hard to interpret whether this variation is due to aging itself or due to cohort-specific influences.
Underlying these debates is a broader methodological issue: distinguishing senescent mortality from deaths caused by external or non-age-related factors. Without a clear separation, as emphasized by Vaupel and Missov, (2014), variations in $b$ are hard to interpret as genuine signals of biological change.
Taken together, the literature suggests that $b$ varies — but none has fully disentangled the sources of that variation. Is it real biological drift? Or is it an illusion — the result of latent period effects, cohort-specific shocks, or observational noise?
To answer this, we move beyond measuring $b$ to explaining it. We decompose its variation into interpretable components: a stable biological rate, a possible cohort trend, and the the accumulated effects of period shocks. This approach allows us to disentangle true biological signals from the noise of history — and to test whether the rate of aging is truly changing, or only appears to.
3 Methodology
Our analysis proceeds in two steps. First, we isolate the portion of mortality attributable to aging — senescent mortality — using a mixture model. Second, we estimate the cohort-specific rate of aging ($b_{t}$) from this senescent component and decompose its variation into interpretable parts.
3.1 Separating Senescent and Non-Senescent Mortality
Total mortality is a blend of different risks: some linked to aging (e.g., degenerative conditions), others associated with external or non-age-related causes (e.g., accidents, infections). To separate these components, we adopt a mixture modeling approach similar to the one proposed by Patricio et al., (2023), which partitions mortality into senescent and non-senescent components.
Formally, let $T$ denote the time of death and $C$ a latent variable indicating cause of death. We assume two broad causes: $c_{1}$ (senescence) and $c_{2}$ (non-senescence). The overall density of $T$ is modeled as a mixture:
$f(x)=\pi_{1}f(x|C=c_{1})+\pi_{2}f(x|C=c_{2}),$ (1)
where $f(x|C=c_{i})$ is the conditional density of age at death given $c_{i}$ and $\pi_{i}=\mathbb{P}(C=c_{i})$, for $i=1,2$.
The senescent component, $f(x|C=c_{1})$, is modeled with a gamma-Gompertz distribution, capturing the exponential increase in mortality risk with age, adjusted for frailty. The non-senescent component, $f(x|C=c_{2})$, is estimated flexibly using penalized splines on the log-hazard scale, with Lasso regularization to prevent overfitting. This setup allows the model to absorb irregularities in early- and mid-life mortality without imposing a strict functional form.
To isolate senescent mortality, we apply the framework described by Patricio and Missov, (2024), which systematically removes non-senescent death counts while preserving the structure and variability of the original mortality data, including its natural variability. This results in a cleaned mortality surface that more closely reflects the biological process of aging.
3.2 Estimating the Rate of Aging, $b$
This mortality surface — denoted $\bar{\mu}(x)$ — represents age-specific mortality rates after the removal of non-senescent deaths. While not a perfect measure of senescence, it substantially reduces non-senescent influence and provides a clearer approximation of aging-related mortality. From this surface, we estimate the cohort-specific Gompertz slope, $b_{t}$, using the transformation introduced by Vaupel, (2022):
$\xi(x)=\log\left(\frac{1}{\bar{\mu}(x)}-\frac{1}{\mu^{*}}\right)\approx\log(\mu_{0})+bx$ (2)
Here, $\mu^{*}$ is the estimated mortality plateau at extreme ages, obtained using the method of Missov and Patricio, (2024). Fitting a straight line to $\xi(x)$ yields $b$, the Gompertz slope — a direct measure of how fast mortality accelerates with age. This method minimizes the influence of early-life mortality and irregularities, allowing for a clean estimation of the aging rate.
3.3 Decomposing Variation in $b$
When we estimate $b$ cohort by cohort, we observe subtle but persistent fluctuations. These movements are too structured to be dismissed as noise — yet they don’t follow a clear trend. Figure 1 illustrates this drift, alongside its first differences and a possible cumulative component.
Figure 1: Top: estimated rate of aging ($\log b$) across French male cohorts. Middle: Cohort-to-cohort differences in $\log b_{t}$, showing cohort-to-cohort shifts. Bottom: possible cumulative period effect. The shape of the drift suggests a common historical influence, motivating the hypothesis that most of the apparent variation in $b$ may reflect period shocks, not biological change.
This pattern raises a central question: what if these changes aren’t biological at all? Period events — such as wars, pandemics, or economic disruptions — affect many cohorts at once, just at different ages. If their effects accumulate over time, they may leave a cohort-shaped fingerprint in the estimates. What appears to be a change in $b$ might simply be the echo of shared historical experience — not a shift in the biology of aging.
To investigate this idea, we model $b_{t}$ as the sum of a constant biological rate, a potential trend, and a cumulative latent period effect. The goal is not only to track variation in aging — but to understand what drives it.
$\log b_{t}=\log b+\beta t+\theta X_{t}+\varepsilon_{t},\quad$ (3)
with
$\quad X_{t}=X_{t-1}+\eta_{t},$ (4)
where
  • $b$ is the constant biological rate of aging,
  • $\beta t$ captures any deterministic trend across birth cohorts,
  • $X_{t}$ is a random walk representing the cumulative effect of period shocks,
  • $\theta$ scales the impact of these shocks on aging,
  • $\varepsilon_{t}$ is cohort-level observational noise,
  • $\eta_{t}$ is white noise driving the latent period process.
This model treats the cumulative period component, $X_{t}$, as a stochastic process: a random walk without drift. Specifically, we assume $\eta_{t}\sim\mathcal{N}(0,\sigma_{X})$ and $\varepsilon_{t}\sim\mathcal{N}(0,\sigma_{b})$, with $\sigma_{X}=1$ fixed to prevent identifiability issues. This choice standardizes the scale of $X_{t}$, allowing us to estimate $\theta$ as the scaling factor that links period shocks to observed changes in aging rates.
The absence of drift in the random walk reflects a conservative assumption: that historical shocks do not push aging rates in a systematically upward or downward direction over time, but instead accumulate in both directions with equal probability. While a drifted process would imply directional historical pressure on aging, our specification treats $X_{t}$ as a mean-zero walk — one that can rise, fall, or reverse, but tends to return toward its earlier levels over time. This aligns with the idea that period shocks are transient in origin, even if their echoes linger across cohorts.
This structure allows us to disentangle what looks like change in $b$ into three interpretable components:
  1. 1.
    A stable biological core ($b$),
  2. 2.
    A potential trend across cohorts ($\beta t$),
  3. 3.
    A latent, cumulative signature of shared historical shocks ($\theta X_{t}$).
By estimating each of these, we can assess whether the observed variation in $b$ reflects true change in the biology of aging — or the subtle accumulation of period-driven distortions over time.
3.4 Estimation Process
We implemented the model in a Bayesian framework using Stan (Stan Development Team, 2025b, ), via its R interface RStan (Stan Development Team, 2025a, ). Four independent Markov chains were run for 6,000 iterations each, with the first 4,000 iterations discarded as warm-up and the remaining 8,000 retained as posterior samples. Convergence was assessed using the $\hat{R}$ diagnostic, which remained below 1.05 for all parameters (Vehtari et al., , 2021), indicating reliable mixing and convergence.
Point estimates reported are the maximum a posteriori (MAP) values, with uncertainty summarized using 95% Highest Posterior Density (HPD) intervals (Patricio and Missov, , 2023).
4 Results
We apply our decomposition model to cohort mortality data from the Human Mortality Database (HMD, , 2025), covering male and female complete birth cohorts after 1850 in France, Denmark, Italy, and Sweden. These results shed light on a central question: does the rate of aging truly change — or merely appear to?
4.1 Observed Variation in the Rate of Aging
Figure 2 shows cohort-specific estimates of the Gompertz slope ($b_{t}$), our measure of the rate of aging, after removing non-senescent mortality. The curves fluctuate — but not wildly. They do not trend upward or downward in a sustained way. They do not jump or break. Instead, they drift: smooth, slow movements across cohorts, too structured to be noise, too subtle to suggest a shift in the biology of aging.
Figure 2: Cohort-specific estimates of $b_{t}$ (rate of aging) across France, Denmark, Italy, and Sweden, by sex. Lines show smooth but structured fluctuations, with no clear long-term trend.
French males offer a striking example. Their estimated rate of aging falls from the 1860 cohort, peaks around 1880, falls slightly, then rises again near 1910. It moves, but it moves like a tide — not a storm. Similar patterns appear elsewhere, especially among males, while females show more muted variation.
Figure 3: Cohort-to-cohort differences in $\log b_{t}$ (i.e., $\log b_{t}-\log b_{t-1}$) across France, Denmark, Italy, and Sweden, by sex. Most values fluctuate gently around zero, with no persistent trend, reinforcing the idea that changes in the estimated rate of aging are smooth and mean-reverting — consistent with the hypothesis of a latent cumulative process such as a random walk.
To further explore the structure of variation in $\log b_{t}$, we plot the first differences across cohorts (Figure 3). These differences cluster closely around zero, with no indication of sustained directional change. The smooth, mean-reverting nature of these fluctuations supports the idea that what we observe is not biological shift, but a slow drift consistent with a latent cumulative process — such as a random walk. This interpretation is further strengthened by formal stationarity tests (ADF, KPSS, Phillips-Perron for $\alpha=0.05$), which confirm that the log-differences are stationary, suggesting $\log b_{t}$ behaves like an integrated process (e.g., Hamilton, , 2020).
This raises the central question: are we seeing biological change? Or are these movements the fingerprint of shared historical events?
4.2 What Drives the Variation?
Decomposition helps us answer that question. Figure 4 isolates the latent cumulative period effect — the component designed to capture shared shocks that drift over time. This effect accounts for the bulk of the observed variation in $\log b_{t}$. It rises, falls, and reverts — not unlike a random walk — and its shape mirrors historical events: wars, epidemics, and social upheavals that left no cohort untouched.
Figure 4: Estimated cumulative period effect ($X_{t}$) by sex and country. These smoothed trajectories explain most of the drift in $\log b_{t}$, supporting the idea that historical shocks, not biological change, drive the variation.
Cohorts carry the echo of past events, even when those events occurred outside their birth year. The fact that this component alone explains most of the structured drift in aging rates offers strong support for the idea that the rate of aging itself may not be changing at all.
4.3 What Remains After the Drift?
The trend component ($\beta t$), shown in Figure 5, is consistently near zero. It captures no systematic rise or fall in $b_{t}$ over birth cohorts. It suggests that, after accounting for period-driven drift, there is no evidence of a directional shift in the pace of aging.
Figure 5: Estimated deterministic trend ($\beta t$) across countries and sexes. The near-zero slopes indicate no cohort-based acceleration or deceleration in aging.
What’s left are the residuals: small, irregular fluctuations shown in Figure 6. They pass standard stationarity tests (ADF, KPSS, Phillips-Perron for $\alpha=0.05$), and they carry no visible signal. In other words, once we account for shared shocks, what remains is consistent with random noise.
Figure 6: Residual variation ($\varepsilon_{t}$) in $\log b_{t}$ after accounting for trend and cumulative shocks. The pattern is stationary and unstructured — consistent with noise, not signal.
4.4 A Built-in Pace
Figure 7 summarizes the estimated decomposition across all countries. The biological rate of aging, $b$, is remarkably consistent and around $0.11$. The deterministic trend, $\beta$, hovers near zero. And while the strength of the period effect, $\theta$, varies modestly, it is always significant — confirming that shared history leaves a real imprint on cohort trajectories.
Figure 7: Posterior estimates of $b$ (left), $\beta$ (center), and $\theta$ (right) by sex and country. The biological rate of aging is stable; the deterministic trend is negligible; and the cumulative effect of period shocks varies in scale but not in kind.
Taken together, these results strengthen Vaupel’s original hypothesis. The pace of aging — as captured by $b$ — appears to be a built-in feature of human biology. The fluctuations we observe are not signals of changing senescence, but echoes of history layered onto data. Aging does not appear to be speeding up or slowing down. It appears to be stable.
5 Discussion
This study set out to revisit Vaupel’s hypothesis: that the rate of aging, $b$, is a biological constant. We asked whether the fluctuations in $b$ observed across cohorts reflect real shifts in the pace of senescence — or whether they arise from something else: the cumulative effect of shared historical shocks.
The results strongly support the latter. Once we isolate senescent mortality and decompose the variation in $b_{t}$, we find no statistical trend. The apparent shifts in the rate of aging are not consistent with a changing biology. Instead, they trace the signature of a latent process that builds slowly over time — a cumulative period effect. In this view, aging itself remains stable. What moves is the historical context.
This reinforces what Vaupel originally proposed: that aging is not slowing down, only starting later. The slope stays the same — it’s just been pushed forward. Our findings align with earlier studies that found limited variation in $b$ (e.g., Barbi et al., , 2003), but offer a new explanation for why those variations appear in the first place. Where previous research saw fluctuation, we show structure.
At the same time, the evidence is not definitive. The confidence intervals around $b_{t}$ — particularly for small population countries — remain wide. The results are statistically consistent with stability, but more precise estimates would require additional data: broader coverage, more cohorts, or longer time series. These would allow for tighter inference and a clearer distinction between signal and noise.
There are also modeling assumptions that should be noted. The latent period effect is structured as a random walk, which captures gradual accumulation but may overlook abrupt shocks or nonlinear interactions. The decomposition is linear, and we treat frailty as stable across cohorts — assumptions that may not hold in all settings. Future work could test alternative specifications, introduce flexible latent processes, or incorporate Bayesian structural time series approaches.
Finally, while this study focuses on national-level mortality, the approach could be extended to explore heterogeneity across subgroups. Do historical shocks leave different fingerprints by sex, education, or socioeconomic status? Does the filtering of senescence vary by social context? These are open questions that matter — not just for understanding aging, but for anticipating how populations will age in the future.
This is, to our knowledge, the first study to explicitly decompose the cohort-level variation in $b$ into a biological constant, a trend, and a latent period effect. In doing so, we offer a framework for biodemographic filtering — a way to separate the signal of aging from the noise of period events. The model is simple, extensible, and grounded in demographic logic. As interest in senescence deepens, we hope this structure helps clarify what, exactly, is changing — and what is not.
6 Conclusion
The Gompertz slope, $b$, is one of the most widely used measures of the pace of aging. Yet its apparent variation across cohorts has raised persistent doubts about whether it is truly stable — or quietly changing.
This study shows that most of the fluctuation in $b$ is not biological. Once we filter out non-senescent mortality and account for the accumulation of shared period effects, $b$ becomes strikingly consistent. There is no meaningful trend, and no evidence that the biology of aging is shifting across cohorts.
These findings provide new support for Vaupel’s hypothesis: that aging proceeds at a fixed biological pace. What has changed over time is not how fast people age, but when aging begins. The rhythm stays the same — it simply starts later.
In the end, what looks like a shift in aging is not a change in senescence, but the echo of history. By separating signal from noise, we gain a clearer view of how aging unfolds — and what, exactly, has changed in the story of human longevity.
Acknowledgments
Silvio C. Patricio gratefully acknowledges the financial support from AXA Research Fund through the funding for the “AXA Chair in Longevity Research”.